# exampledefaultouterlinewidth=0pt,innerlinewidth=0pt, roundcorner=5pt,backgroundcolor=myblue, leftmargin=0.2cm,rightmargin=0.2cm

### v1

@@ Abstract
We develop a framework for analyzing the training and learning rate dynamics on a large class of high-dimensional optimization problems, which we call the high line, trained using one-pass stochastic gradient descent (SGD) with adaptive learning rates. We give exact expressions for the risk and learning rate curves in terms of a deterministic solution to a system of ODEs. We then investigate in detail two adaptive learning rates – an idealized exact line search and AdaGrad-Norm – on the least squares problem. When the data covariance matrix has strictly positive eigenvalues, this idealized exact line search strategy can exhibit arbitrarily slower convergence when compared to the optimal fixed learning rate with SGD. Moreover we exactly characterize the limiting learning rate (as time goes to infinity) for line search in the setting where the data covariance has only two distinct eigenvalues. For noiseless targets, we further demonstrate that the AdaGrad-Norm learning rate converges to a deterministic constant inversely proportional to the average eigenvalue of the data covariance matrix, and identify a phase transition when the covariance density of eigenvalues follows a power law distribution.
@@ 1 Introduction
There are many adaptive learning rate strategies used in machine learning with many design goals. Some are known to adapt to SGD gradient noise while others are robust to hyper-parameters (e.g., [59, 4]). Theoretical results for adaptive algorithms tend to focus on guaranteeing minimax-optimal rates, but this theory is not engineered to provide realistic performance comparisons; indeed many adaptive algorithms are minimax-optimal, and so more precise statements are needed to distinguish them. For instance, the exact learning rates (or rate schedules) to which these strategies converge are unknown, nor their dependence on the geometry of the problem. Moreover, we often do not know how these adaptive stepsizes compare with well-tuned constant or decaying fixed learning rate stochastic gradient descent (SGD), which can be viewed as a cost associated with selecting the adaptive strategy in comparison to tuning by hand.
@@
Figure 1: Concentration of learning rate and risk for AdaGrad-Norm on least squares with label noise $\omega=1$ (left) and logistic regression with no noise (right). As dimension increases, both risk and learning rate concentrate around a deterministic limit (red) described by our ODE in Theorem 2.1. The initial risk increase (left) suggests the learning rate started too high, but AdaGrad-Norm adapts. Our ODEs predict this behavior. See Sec.  12 for simulation details.
We suppose $a\sim\mathcal{N}(0,K)$ where $K\in\mathbb{R}^{d\times d}$ is the covariance matrix. We train (1) using (one-pass) stochastic gradient descent with adaptive learning rates, $\mathfrak{g}_{k}$ (SGD+AL). Our main goal is to give a framework for better [^1] More realistic, in that it deals with high-dimensional anisotropic loss geometries and more precise, in that it can distinguish minimax optimal algorithms as more-or-less performant. performance analysis of these adaptive methods. We then illustrate this framework by considering two adaptive learning rate algorithms on the least squares problem [^2] We extend some results to the general strongly convex setting., the results of which appear in Table LABEL:table:learning_rates: exact line-search (idealistic) (Sec. 3) and AdaGrad-Norm (Sec. 4). We fully expect other losses and adaptive learning rates can be studied using this approach.
@@
Greed can be arbitrarily bad in the presence of strong anisotropy. Our analysis reveals that exact line search, which is to say optimally decreasing the risk at each step, can run arbitrarily slower than the best fixed learning rate for SGD on a least squares problem when $\lambda_{\min}\stackrel{{\scriptstyle\text{def}}}{{=}}\lambda_{\min}(K)>C>0$. The best fixed stepsize (least squares problem) is $(\text{Tr}(K)/d)^{-1}$ or the inverse of the average eigenvalue, see Polyak stepsize [46]. Line search, on the other hand, converges to a fixed stepsize of order $\lambda_{\min}/(\text{Tr}(K^{2})/d)$. It can be that $\lambda_{\min}/(\text{Tr}(K^{2})/d)\ll(\text{Tr}(K)/d)^{-1}$ making exact line search substantially underperform Polyak stepsize. We further explore this and, in the case where $d$-eigenvalues of $K$ take only two values $\lambda_{1}>\lambda_{2}>0$, we give an exact expression as a function of $\lambda_{1}$ and $\lambda_{2}$ for the limiting behavior of $\gamma_{t}$ as $t\to\infty$ (See Fig. 5).
@@
Notation.
Define $\mathbb{R}_{+}=[0,\infty)$. We say an event holds with overwhelming probability, w.o.p., if there is a function $\omega:\mathbb{N}\to\mathbb{R}$ with $\omega(d)/\log d\to\infty$ so that the event holds with probability at least $1-e^{-\omega(d)}$. We let $\bm{1}_{A}(x)$ be the indicator function of the set $A$ where it is $1$ if $x\in A$ and 00 otherwise. For a matrix $A\in\mathbb{R}^{m\times d}$, we use $\|A\|$ to denote the Frobenius norm and $\|A\|_{\text{op}}$ to denote the operator-2 norm. If unspecified, we assume that the norm is the Frobenius norm. For normed vector spaces $\mathcal{A}$, $\mathcal{B}$ with norms $\|\cdot\|_{\mathcal{A}}$ and $\|\cdot\|_{\mathcal{B}}$, respectively, and for $\alpha\geq 0$, we say a function $F\,:\,\mathcal{A}\to\mathcal{B}$ is $\alpha$-pseudo-Lipschitz with constant $L$ if for any $A,\hat{A}\in\mathcal{A}$, we have
@@
AdaGrad-Norm.
AdaGrad, introduced by [18, 35], updates the learning rate at each iteration using the stochastic gradient information. The single stepsize version [55, 31, 57], that depends on the norm of the gradient, (see Table LABEL:table:learning_rate_expressions for the updates), has been shown to be robust to input parameters [33]. Several works have shown worst case convergence guarantees [55, 32, 20, 53]. A linear rate of $O(\exp(-\kappa T))$ is possible for $\mu$-strongly convex, $L$-smooth functions ($\kappa$ is the condition number $\mu/L$). In [58] (similar idea in [57]), the authors show for strongly-convex, smooth stochastic objectives (with additional assumptions) that the AdaGrad-Norm learning rate exhibits a two stage behavior – a burn in phase and then when it reaches the smoothness constant it self-stablizes.
Stochastic line search and Polyak stepsizes.
Recently there has been renewed interest in studying stochastic line search [41, 51, 19] and stochastic Polyak stepsize (and their variants) [34, 29, 40, 7, 25, 47, 38, 26]. Much of this research focuses on worst-case convergence guarantees for strongly convex and smooth functions (see e.g., [34]) and designing practical algorithms. In [50], the authors provide a bound on the learning rate for Armijo line search in the finite sum setting with a rate of $L_{\max}/\text{avg. $\mu$}$ where avg. $\mu$ is the avg. strong convexity and $L_{\max}$ is the max. Lipschitz constant of the individual functions. In this work, we consider a slightly different problem. We work with the population loss and we note that the analogue to $L_{\max}$ for us would require that the samples $a$ satisfy $\|aa^{T}\|_{\text{op}}\leq L_{\max}$ for all $a$; this fails to hold for $a\sim\mathcal{N}(0,K)$. Moreover, $L_{\max}$ could be much worse than ${\mathbb{E}}\,[\|aa^{T}\|_{\text{op}}]$.
Deterministic dynamics of stochastic algorithms in high-dimensions.
The literature on deterministic dynamics for isotropic Gaussian data has a long history [48, 8, 9, 49]. These results have been rigorously proven and extended to other models under the isotropic Gaussian assumption [22, 54, 2, 6, 15, 1, 16]. Extensions to multi-pass SGD with small mini-batches [44] as well as momentum [30] have also been studied. Other high-dimensional limits leading to a different class of dynamics also exist [36, 21, 11, 12, 10]. Recently, significant contributions have been made in understanding the effects of a non-identity data covariance matrix on the training dynamics [13, 5, 24, 60, 23, 14]. The non-identity covariance modifies the optimization landscape and affects convergence properties [14]. This work extends [14] to stochastic adaptive algorithms, exploring the effect of non-identity covariance in these algorithms.
@@ 1.1 Model Set-up
For $a,X,X^{\star}\in\mathbb{R}^{d}$, $\epsilon\in\mathbb{R}$, and a function $f:\mathbb{R}^{3}\rightarrow\mathbb{R}$, we seek to minimize an expected risk function $\mathcal{R}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ of the form
$\mathcal{R}(X)\stackrel{{\scriptstyle\text{def}}}{{=}}{\mathbb{E}}\,_{a,\epsilon}[\Psi(X;a,\epsilon)]\quad\text{for}\quad(a,\epsilon)\sim\mathcal{D},\quad\text{and}\quad\Psi(X;a,\epsilon)=f(\langle a,X\rangle,\langle a,X^{\star}\rangle,\epsilon),$ (2)
which we call the high-dimensional linear composite [^3] Note that $d$ need not be large to define this, but the structure allows us to consider $d$ as a tunable parameter. Moreover, as we increase $d$, the analysis we do will be more meaningful.. In what follows, we use the matrix $W=[X|X^{\star}]\in\mathbb{R}^{d\times 2}$ that concatenates $X$ and $X^{\star}$, and we shall let $B=B(W)=W^{T}KW$ be the covariance matrix of the Gaussian vector $(\langle{a,X}\rangle,\langle{a,X^{\star}}\rangle)$. We assume on $f$.
@@
By assumption, $\mathcal{R}(X)$ involves an expectation over the correlated Gaussians $\langle{a,X}\rangle$ and $\langle{a,X^{\star}}\rangle$. We may represent this expectation $\mathcal{R}(X)\stackrel{{\scriptstyle\text{def}}}{{=}}h(B)$ for some function $h\,:\,\mathbb{R}^{2\times 2}\to\mathbb{R}$. We shall require that this function $h$ is well-behaved.
@@
The final assumption is the well-behavior of the Fisher information matrix of the gradients. The first coordinate of $f$ is special, in that the optimizer must be able to differentiate it, and hence it is convenient to consider $f(x,x^{\star},\epsilon)$ as a function of a single variable with two parameters, i.e. $f(x,x^{\star},\epsilon)=f(x;x^{\star},\epsilon)$, and so we write $f^{\prime}$ for the (almost everywhere) derivative of $f$ in its first variable.
@@
A large class of natural regression problems fit within this framework, such as logistic regression and least squares (see [14, Appendix B]). We also note that Assumptions 3 and 4 are nearly satisfied for $L$-smooth objectives $f$ (see Lemma 6.1), and a version of the main theorem holds under just this assumption (albeit with a weaker conclusion).
@@ 1.2 Algorithmic set-up
where $\mathfrak{g}_{k}>0$ is a learning rate (see assumptions below). To perform our analysis, we place the following assumption on the initialization $X_{0}$ and signal $X^{\star}$.
@@
The learning rate function $\gamma\,:\,\mathbb{R}_{+}\times D([0,\infty))\times D([0,\infty))\times D([0,\infty))\to\mathbb{R}$ is $\alpha$-pseudo-Lipschitz with constant $L(\gamma)$ (independent of $d$) in $D([0,\infty))\times D([0,\infty))\times D([0,\infty))$. Moreover for some constant $C=C(\gamma)>0$ independent of $d$ and $\delta>0$,
@@
The inequality (5) ensures that the learning rate concentrates around the mean behavior of the stochastic gradients. Many well-known adaptive stepsizes satisfy (4) and Assumption 6 including AdaGrad-Norm, DoG, D-Adaptation, and RMSProp (see Table LABEL:table:learning_rate_expressions, Sec. 5, and Sec. 7.3).
@@ 2 Deterministic dynamics for SGD with adaptive learning rates
The risk $\mathcal{R}(X)$ and Fisher matrix can be evaluated solely in terms of the covariance matrix $B$. Thus, to know the evolution of the risk over time, it would suffice to know the evolution of $B$. Alas, except in the isotropic case where $K$ is a multiple of the identity, the evolution of $B$ is not autonomous (i.e., its time evolution depends on other unknown variables). However, if we let $(\lambda_{i},\omega_{i})$ be the eigenvalues and corresponding orthonormal eigenvectors of $K$, we can consider projections $V_{i}(X_{k})=d\cdot W_{k}^{T}\omega_{i}\omega_{i}^{T}W_{k}$, and it turns out that these behave autonomously. Formally we do this below, building on [14], in which this was done with non-adapted learning rates.
@@
This time change is necessary, as when we scale the size of the problem, more time is needed to solve the underlying problem. This scaling law scales SGD so all training dynamics live on the same space. One can solve a smaller $d$ problem and scale it to recover the training dynamics of the larger problem.
@@
$\mathrm{d}\mathscr{V}_{11,i}(t)$ $=-2\lambda_{i}\gamma_{t}\left(\mathscr{V}_{11,i}(t)H_{1,t}+H_{1,t}\mathscr{V}_{11,i}(t)+\mathscr{V}_{12,i}(t)H_{2,t}\right)+\lambda_{i}\gamma_{t}^{2}\mathscr{I}(t),$ (7)
@@
In fact, we can derive deterministic dynamics for a large class of statistics which are linear combinations of $\mathscr{V}(t)$ and functions thereof (See Theorem 6.1, and Corollary 6.1).
@@
Example: Least Squares.
One canonical example of (2) is least squares where we aim to recover the target $X^{\star}$ given noisy observations $\langle a,X^{\star}\rangle+\epsilon$. In this case, the least squares problem is
$\mathscr{R}(t)$ $=\tfrac{1}{2}(X_{0}-X^{\star})^{T}Ke^{-2K\int_{0}^{t}\gamma_{s}\,\mathop{}\!\mathrm{d}s}(X_{0}-X^{\star})+\tfrac{1}{2}\omega^{2}$ (11)
$\qquad+\tfrac{1}{d}\textstyle\int_{0}^{t}\gamma_{s}^{2}\text{Tr}(K^{2}e^{-2K\int_{s}^{t}\gamma_{\tau}\mathop{}\!\mathrm{d}\tau})\mathscr{R}(s)\,\mathop{}\!\mathrm{d}s.$
This is a convolution Volterra equation, and it has appeared earlier in [13, 45, 44, 43]. The convergence threshold of this equation is simply $\gamma_{t}<\tfrac{2d}{\text{Tr}K}$.
Recall $((\lambda_{i},\omega_{i}):1\leq i\leq d)$ are the eigenvalues and orthonormal eigenvectors of $K$. In the noiseless label case (i.e., $\epsilon=0$), the risk $\mathscr{R}(t)=\tfrac{1}{2d}\sum_{i=1}^{d}\lambda_{i}\mathscr{D}_{i}^{2}(t)$. Using the ODEs in (7), we get the following deterministic equivalent ODE for the $\mathscr{D}^{2}_{i}$’s:
We now perform a deep analysis of the dynamics of the learning rate on least squares (10); this analysis generalizes to a setting where the outer function $f$ is strongly convex (see 8.1).
Figure 2: Comparison for Exact Line Search and Polyak Stepsize on a noiseless least squares problem. The left plot illustrates the convergence of the risk function, while the right plot depicts the convergence of the quotient $\gamma_{t}/\frac{\lambda_{\min}(K)}{\frac{1}{d}\operatorname{Tr}(K^{2})}$ for Polyak stepsize and exact line search. Both plots highlight the implication of equation (14) in high-dimensional settings, where a broader spectrum of $K$ results in $\frac{\lambda_{\min}(K)}{\tfrac{1}{d}\mathrm{Tr}(K^{2})}\ll\frac{1}{\tfrac{1}{d}\mathrm{Tr}(K)}$, indicating slower risk convergence and poorer performance of exact line search (unmarked) as it deviates from the Polyak stepsize (circle markers) . The gray shaded region demonstrates that equation (14) is satisfied. See Appendix 12 for simulation details.

### v2

@@ Abstract
We develop a framework for analyzing the training and learning rate dynamics on a large class of high-dimensional optimization problems, which we call the high line, trained using one-pass stochastic gradient descent (SGD) with adaptive learning rates. We give exact expressions for the risk and learning rate curves in terms of a deterministic solution to a system of ODEs. We then investigate in detail two adaptive learning rates – an idealized exact line search and AdaGrad-Norm – on the least squares problem. When the data covariance matrix has strictly positive eigenvalues, this idealized exact line search strategy can exhibit arbitrarily slower convergence when compared to the optimal fixed learning rate with SGD. Moreover we exactly characterize the limiting learning rate (as time goes to infinity) for line search in the setting where the data covariance has only two distinct eigenvalues. For noiseless targets, we further demonstrate that the AdaGrad-Norm learning rate converges to a deterministic constant inversely proportional to the average eigenvalue of the data covariance matrix, and identify a phase transition when the covariance density of eigenvalues follows a power law distribution. We provide our code for evaluation at https://github.com/amackenzie1/highline2024.
@@ 1 Introduction
There are many adaptive learning rate strategies used in machine learning with many design goals. Some are known to adapt to stochastic gradient descent (SGD) gradient noise while others are robust to hyper-parameters (e.g., [63, 4]). Theoretical results for adaptive algorithms tend to focus on guaranteeing minimax-optimal rates, but this theory is not engineered to provide realistic performance comparisons; indeed many adaptive algorithms are minimax-optimal, and so more precise statements are needed to distinguish them. For instance, the exact learning rates (or rate schedules) to which these strategies converge are unknown, nor their dependence on the geometry of the problem. Moreover, we often do not know how these adaptive stepsizes compare with well-tuned constant or decaying fixed learning rate SGD, which can be viewed as a cost associated with selecting the adaptive strategy in comparison to tuning by hand.
@@
Figure 1: Concentration of learning rate and risk for AdaGrad-Norm on least squares with label noise $\omega=1$ (left) and logistic regression with no noise (right). As dimension increases, both risk and learning rate concentrate around a deterministic limit (red) described by our ODE in Theorem 2.1. The initial risk increase (left) suggests the learning rate started too high, but AdaGrad-Norm adapts. Our ODEs predict this behavior. See Sec. H for simulation details.
We suppose $a\sim\mathcal{N}(0,K)$ where $K\in\mathbb{R}^{d\times d}$ is the covariance matrix. We train (1) using (one-pass) stochastic gradient descent with adaptive learning rates, $\mathfrak{g}_{k}$ (SGD+AL). Our main goal is to give a framework for better [^1] More realistic, in that it deals with high-dimensional anisotropic loss geometries and more precise, in that it can distinguish minimax optimal algorithms as more-or-less performant. performance analysis of these adaptive methods. We then illustrate this framework by considering two adaptive learning rate algorithms on the least squares problem [^2] We extend some results to the general strongly convex setting., the results of which appear in Table LABEL:table:learning_rates: exact line-search (idealistic) (Sec. 3) and AdaGrad-Norm (Sec. 4). We expect other losses and adaptive learning rates can be studied using this approach.
@@
Greed can be arbitrarily bad in the presence of strong anisotropy (that is, $\text{Tr}(K)/d\ll\text{Tr}(K^{2})/d$). Our analysis reveals that exact line search, which is to say optimally decreasing the risk at each step, can run arbitrarily slower than the best fixed learning rate for SGD on a least squares problem when $\lambda_{\min}\stackrel{{\scriptstyle\text{def}}}{{=}}\lambda_{\min}(K)>C>0$. The best fixed stepsize (least squares problem) is $(\text{Tr}(K)/d)^{-1}$ or the inverse of the average eigenvalue, see Polyak stepsize [48]. Line search, on the other hand, converges to a fixed stepsize of order $\lambda_{\min}/(\text{Tr}(K^{2})/d)$. It can be that $\lambda_{\min}/(\text{Tr}(K^{2})/d)\ll(\text{Tr}(K)/d)^{-1}$ making exact line search substantially underperform Polyak stepsize. We further explore this and, in the case where $d$-eigenvalues of $K$ take only two values $\lambda_{1}>\lambda_{2}>0$, we give an exact expression as a function of $\lambda_{1}$ and $\lambda_{2}$ for the limiting behavior of $\gamma_{t}$ as $t\to\infty$ (See Fig. 5).
@@
Notation. Define $\mathbb{R}_{+}=[0,\infty)$. We say an event holds with overwhelming probability, w.o.p., if there is a function $\omega:\mathbb{N}\to\mathbb{R}$ with $\omega(d)/\log d\to\infty$ so that the event holds with probability at least $1-e^{-\omega(d)}$. We let $\bm{1}_{A}(x)$ be the indicator function of the set $A$ where it is $1$ if $x\in A$ and 00 otherwise. For a matrix $A\in\mathbb{R}^{m\times d}$, we use $\|A\|$ to denote the Frobenius norm and $\|A\|_{\text{op}}$ to denote the operator-2 norm. If unspecified, we assume that the norm is the Frobenius norm. For normed vector spaces $\mathcal{A}$, $\mathcal{B}$ with norms $\|\cdot\|_{\mathcal{A}}$ and $\|\cdot\|_{\mathcal{B}}$, respectively, and for $\alpha\geq 0$, we say a function $F\,:\,\mathcal{A}\to\mathcal{B}$ is $\alpha$-pseudo-Lipschitz with constant $L$ if for any $A,\hat{A}\in\mathcal{A}$, we have
@@
AdaGrad-Norm. AdaGrad, introduced by [19, 36], updates the learning rate at each iteration using the stochastic gradient information. The single stepsize version [59, 32, 61], that depends on the norm of the gradient, (see Table LABEL:table:learning_rate_expressions for the updates), has been shown to be robust to input parameters [34]. Several works have shown worst-case convergence guarantees [59, 33, 21, 57]. A linear rate of $O(\exp(-\kappa T))$ is possible for $\mu$-strongly convex, $L$-smooth functions ($\kappa$ is the condition number $\mu/L$). In [62] (similar idea in [61]), the authors show for strongly convex, smooth stochastic objectives (with additional assumptions) that the AdaGrad-Norm learning rate exhibits a two stage behavior – a burn in phase and then when it reaches the smoothness constant it self-stablizes.
Stochastic line search and Polyak stepsizes. Recently there has been renewed interest in studying stochastic line search [42, 54, 20] and stochastic Polyak stepsize (and their variants) [35, 30, 41, 7, 26, 49, 39, 27]. Much of this research focuses on worst-case convergence guarantees for strongly convex and smooth functions (see e.g., [35]) and designing practical algorithms. In [53], the authors provide a bound on the learning rate for Armijo line search in the finite sum setting with a rate of $L_{\max}/\text{avg. $\mu$}$ where avg. $\mu$ is the avg. strong convexity and $L_{\max}$ is the max. Lipschitz constant of the individual functions. In this work, we consider a slightly different problem. We work with the population loss and we note that the analogue to $L_{\max}$ for us would require that the samples $a$ satisfy $\|aa^{T}\|_{\text{op}}\leq L_{\max}$ for all $a$; this fails to hold for $a\sim\mathcal{N}(0,K)$. Moreover, $L_{\max}$ could be much worse than ${\mathbb{E}}\,[\|aa^{T}\|_{\text{op}}]$.
Deterministic dynamics of stochastic algorithms in high-dimensions. The literature on deterministic dynamics for isotropic Gaussian data has a long history [50, 9, 10, 51]. These results have been rigorously proven and extended to other models under the isotropic Gaussian assumption [23, 58, 2, 6, 16, 1, 17]. Extensions to multi-pass SGD with small mini-batches [46] as well as momentum [31] have also been studied. Other high-dimensional limits leading to a different class of dynamics also exist [37, 22, 12, 13, 11]. Recently, significant contributions have been made in understanding the effects of a non-identity data covariance matrix on the training dynamics [14, 5, 25, 64, 24, 15]. The non-identity covariance modifies the optimization landscape and affects convergence properties, as discussed in [15]. This work extends the findings of [15] to stochastic adaptive algorithms, exploring the effect of non-identity covariance within these algorithms. Notably, Theorem 1.1 from [15] is restricted to deterministic learning rate schedules, limiting its applicability in many practical scenarios. In contrast, our Theorem 2.1 accommodates stochastic adaptive learning rates, aligning with widely used algorithms in practice.
@@ 1.1 Model Set-up
For $a,X,X^{\star}\in\mathbb{R}^{d}$, $\epsilon\in\mathbb{R}$, and a function $f:\mathbb{R}^{3}\rightarrow\mathbb{R}$, we seek to minimize an expected risk function $\mathcal{R}:\mathbb{R}^{d}\rightarrow\mathbb{R}$, which we refer to as the high-dimensional linear composite [^3] Note that $d$ need not be large to define this, but the structure allows us to consider $d$ as a tunable parameter. Moreover, as we increase $d$, the analysis we do will be more meaningful., of the form
$\mathcal{R}(X)\stackrel{{\scriptstyle\text{def}}}{{=}}{\mathbb{E}}\,_{a,\epsilon}[\Psi(X;a,\epsilon)]\quad\text{for}\quad(a,\epsilon)\sim\mathcal{D},\quad\text{and}\quad\Psi(X;a,\epsilon)=f(\langle a,X\rangle,\langle a,X^{\star}\rangle,\epsilon).$ (2)
In what follows, we use the matrix $W=[X|X^{\star}]\in\mathbb{R}^{d\times 2}$ that concatenates $X$ and $X^{\star}$, and we shall let $B=B(W)=W^{T}KW$ be the covariance matrix of the Gaussian vector $(\langle{a,X}\rangle,\langle{a,X^{\star}}\rangle)$.
@@
By assumption, $\mathcal{R}(X)$ involves an expectation over the correlated Gaussians $\langle{a,X}\rangle$ and $\langle{a,X^{\star}}\rangle$. We can express this as $\mathcal{R}(X)\stackrel{{\scriptstyle\text{def}}}{{=}}h(B)$ for some well-behaved function $h\,:\,\mathbb{R}^{2\times 2}\to\mathbb{R}$.
@@
The final assumption is the well-behavior of the Fisher information matrix of the gradients. The first coordinate of $f$ is special, as the optimizer must be able to differentiate it. Thus, we treat $f(x,x^{\star},\epsilon)$ as a function of a single variable with two parameters: $f(x,x^{\star},\epsilon)=f(x;x^{\star},\epsilon)$ and denote the (almost everywhere) derivative with respect to the first variable as $f^{\prime}$.
@@
A large class of natural regression problems fit within this framework, such as logistic regression and least squares (see [15, Appendix B]). We also note that Assumptions 3 and 4 are nearly satisfied for $L$-smooth objectives $f$ (see Lemma B.1), and a version of the main theorem holds under just this assumption (albeit with a weaker conclusion).
@@ 1.2 Algorithmic set-up
where $\mathfrak{g}_{k}>0$ is a learning rate (see assumptions below) [^4] Note that cases where $\frac{\operatorname{Tr}(K^{2})}{d}=o(d)$ can lead to dynamics that converge to full-batch gradient flow. While our theorem specifically addresses the scenario where the intrinsic dimension, $\text{Dim}(K)\stackrel{{\scriptstyle\text{def}}}{{=}}\operatorname{Tr}(K)/\|K\|_{\text{op}}$, satisfies $\text{Dim}(K)=\Theta(d)$, other cases, such as $\text{Dim}(K)=o(d)$, may require different learning rate scalings.. To perform our analysis, we place the following assumption on the initialization $X_{0}$ and signal $X^{\star}$.
@@
The learning rate function $\gamma\,:\,\mathbb{R}_{+}\times D([0,\infty))\times D([0,\infty))\times D([0,\infty))\to\mathbb{R}$ is $\alpha$-pseudo-Lipschitz with constant $L(\gamma)$ (independent of $d$) in $D([0,\infty))\times D([0,\infty))\times D([0,\infty))$. Moreover, for some constant $C=C(\gamma)>0$ independent of $d$ and $\delta>0$,
@@
The inequality (5) ensures that the learning rate concentrates around the mean behavior of the stochastic gradients. Many well-known adaptive stepsizes satisfy (4) and Assumption 6 including AdaGrad-Norm, DoG, D-Adaptation, and RMSProp (see Table LABEL:table:learning_rate_expressions, Sec. A, and Sec. C.3).
@@ 2 Deterministic dynamics for SGD with adaptive learning rates
The risk $\mathcal{R}(X)$ and Fisher matrix can be evaluated solely in terms of the covariance matrix $B$. Thus, to know the evolution of the risk over time, it would suffice to know the evolution of $B$. Alas, except in the isotropic case where $K$ is a multiple of the identity, the evolution of $B$ is not autonomous (i.e., its time evolution depends on other unknown variables). However, if we let $(\lambda_{i},\omega_{i})$ be the eigenvalues and corresponding orthonormal eigenvectors of $K$, we can consider projections $V_{i}(X_{k})=d\cdot W_{k}^{T}\omega_{i}\omega_{i}^{T}W_{k}$, and it turns out that these behave autonomously.
Example: Least Squares.
One canonical example of (2) is least squares, where we aim to recover the target $X^{\star}$ given noisy observations $\langle a,X^{\star}\rangle+\epsilon$. In this case, the least squares problem is
$\mathscr{R}(t)=\tfrac{1}{2}(X_{0}-X^{\star})^{T}Ke^{-2K\int_{0}^{t}\gamma_{s}\,\mathop{}\!\mathrm{d}s}(X_{0}-X^{\star})+\tfrac{1}{2}\omega^{2}+\tfrac{1}{d}\textstyle\int_{0}^{t}\gamma_{s}^{2}\text{Tr}(K^{2}e^{-2K\int_{s}^{t}\gamma_{\tau}\mathop{}\!\mathrm{d}\tau})\mathscr{R}(s)\,\mathop{}\!\mathrm{d}s.$
This is a convolution Volterra equation with a convergence threshold of $\gamma_{t}<\tfrac{2d}{\text{Tr}K}$ [14, 47, 46, 44].
In the noiseless label case (i.e., $\epsilon=0$), the risk is given by $\mathscr{R}(t)=\tfrac{1}{2d}\sum_{i=1}^{d}\lambda_{i}\mathscr{D}_{i}^{2}(t)$. Using the ODEs in (9), we get the following deterministic equivalent ODE for the $\mathscr{D}^{2}_{i}$’s:
We will perform a deep analysis of the dynamics of the learning rate on least squares (7), which will generalize to settings where the outer function $f$ is strongly convex (see D.1).
@@
This time change is necessary, as when we scale the size of the problem, more time is needed to solve the underlying problem. This scaling law scales SGD so all training dynamics live on the same space. One can solve a smaller $d$ problem and scale it to recover the training dynamics of the larger problem. [^6] Note that, holding time fixed, we perform $O(d)$ gradient updates for a problem of dimension $d$. For the problems considered here, this scaling leads to consistent dynamics, but there do exist related problems where a different scaling is more appropriate. For example, under random initialization, to capture the escape of phase retrieval from the high-dimensional saddle, $O(d\log d)$ iterations are needed; see for example [56].
@@
$\mathrm{d}\mathscr{V}_{11,i}(t)$ $=-2\lambda_{i}\gamma_{t}\left(\mathscr{V}_{11,i}(t)H_{1,t}+H_{1,t}\mathscr{V}_{11,i}(t)+\mathscr{V}_{12,i}(t)H_{2,t}+H_{2,t}\mathscr{V}_{12,i}(t)\right)+\lambda_{i}\gamma_{t}^{2}\mathscr{I}(t),$ (9)
@@
In fact, we can derive deterministic dynamics for a large class of statistics which are linear combinations of $\mathscr{V}(t)$ and functions thereof (See Theorem B.1, and Corollary B.1).
@@
Figure 2: Comparison for Exact Line Search and Polyak Stepsize on a noiseless least squares problem. The left plot illustrates the convergence of the risk function, while the right plot depicts the convergence of the quotient $\gamma_{t}/\frac{\lambda_{\min}(K)}{\frac{1}{d}\operatorname{Tr}(K^{2})}$ for Polyak stepsize and exact line search. Both plots highlight the implication of equation (13) in high-dimensional settings, where a broader spectrum of $K$ results in $\frac{\lambda_{\min}(K)}{\tfrac{1}{d}\mathrm{Tr}(K^{2})}\ll\frac{1}{\tfrac{1}{d}\mathrm{Tr}(K)}$, indicating slower risk convergence and poorer performance of exact line search (unmarked) as it deviates from the Polyak stepsize (circle markers) . The gray shaded region demonstrates that equation (13) is satisfied. See Appendix H for simulation details.